ID |
Date |
Author |
Subject |
Text |
 |
666
|
Sun Apr 6 14:21:53 2025 |
TD | [How To] Startup CARME |
- water flow/interlock checks
check flow
|
|
665
|
Sun Apr 6 11:40:48 2025 |
TD | Options DB reconstructed | Options directory corrupted unable to change
Options settings per https://elog.ph.ed.ac.uk/CARME/664
|
|
664
|
Sun Apr 6 10:17:49 2025 |
TD | Sunday 6 April |
11.15 CARME IE514 ion gauge remains OFF
ML says there are 2x ion gauges with |
15x |
663
|
Sat Apr 5 16:15:04 2025 |
TD | FEE64 power interlock installed |
Water flow and dew point sensors have been
installed on the outlet and inlet cooling |
10x |
662
|
Sat Apr 5 16:13:32 2025 |
TD | Saturday 5 April |
12.00 CARME IE514 ion gauge 4.2e-11 mbar
CARME IE514 OFF
|
7x |
661
|
Fri Apr 4 10:17:16 2025 |
TD | Friday 4 April |
10.43 Reboot carme-gsi - unable to open a
working terminal
|
|
660
|
Mon Mar 10 15:47:27 2025 |
TD | Analysis data files G22-201 R79_10 & R79_24 |
Analysis of data files G22-201 R79_10 and
R79_24 - attachments 1-2
|
 |
659
|
Tue Jul 2 17:43:32 2024 |
TD | Offline analysis G22-201/R89 | Offline analysis data file G22-201/R89_0 to
R89_13 (c. 6h40m data)
|
9x |
658
|
Wed Jun 19 12:11:22 2024 |
CB, RSS | CARME status | DSSD off
FEEs off
Motors out, pin inserted, compressed |
|
657
|
Wed Jun 19 12:08:34 2024 |
CB + various | Run 101-105 | Run 101: no target. See ELOG entry by TD.
Run 102: ??
Run 103: ?? |
|
656
|
Wed Jun 19 12:06:50 2024 |
CB, OF, SS | Run 98/99/100 - higher energy | 10:21 June 17
New energy: 1.125 MeV/u
Target in. Run 98 started. |
|
655
|
Wed Jun 19 12:04:02 2024 |
CB + OF, SS | Run 96/97 | Run 96 - as run 89, while we wait for beam
energy change at higher energies
Run stopped around 08:56 |
|
654
|
Wed Jun 19 12:01:29 2024 |
TD | 23:00-07:00 Wednesday 19 June | 23.18 DSSSD bias & leakage current OK - attachment
1
FEE64 temps OK - attachment 2
|
23x |
653
|
Wed Jun 19 12:00:56 2024 |
CB | Run 93/94/95 - background | 04:00 17 June
Moved the detectors back to the
usual configuration (as per run 89). |
6x |
652
|
Wed Jun 19 11:54:26 2024 |
CB | Run 91/92 - moving detectors | ~ 2:00 17 June
Run 91 |
    |
651
|
Mon Jun 17 23:10:08 2024 |
CB | Run 90 | 00:10
As run 89, detectors moved fully
out. |
 |
650
|
Mon Jun 17 16:26:00 2024 |
JM, RSS, CB | Run 89 | 17:20
(Run 87 Junk and Run 88 does not
exist) |
19x |
649
|
Mon Jun 17 06:26:45 2024 |
JM, OF, AK, CB | Change to 5+ beam | 07:00
We are changing to 5+ beam so that
we can access higher energies and possibly |
7x |
648
|
Mon Jun 17 05:47:35 2024 |
JM | Run 86 | 06:50
No sign of the resonance during
night shift. New run started but when operators |
|
647
|
Mon Jun 17 04:01:16 2024 |
JM | Run 85 | 05:00
Cooler voltage setpoint 247.45
V. |
|